OpenStudy (anonymous):

if f(x) = 10x -4 find f(-5)

OpenStudy (stefrheart):

You need to substitute x with -5 so =10(-5)-4 then you need to multiply =-50- -4 and the rule with subtracting fractions is to leave, change, opposite =-50+4 so the equation equals =-46

OpenStudy (stefrheart):

Do you understand?